# HiReview: Hierarchical Taxonomy-Driven Automatic Literature Review Generation

### v1

@@ 4.2.1 Hierarchical Citation Graph Clustering
This density assesses how densely connected each node is within its local neighborhood. High-density nodes are more likely to be in the core of a cluster, whereas low-density nodes are more likely to be in ambiguous regions between clusters. $\hat{d}_{u}=0$ if $u$ is isolated. Given $p_{uv}$, $\hat{d}_{u}$, $\hat{d}_{v}$, and a pre-defined edge connection threshold $p_{\tau}$, we update the candidate cluster as,
$c_{u}=\{u\}+\{v~{}|~{}\hat{d}_{u}<\hat{d}_{v}~{}\text{and}~{}p_{uv}>p_{\tau}\},$ (4)
where nodes with higher density are added to multiple candidate clusters. This approach avoids prematurely merging large clusters without clear boundaries. Early merging can obscure the nuanced relationships between papers, especially in citation networks where subtopics often overlap. At the first level, a soft clustering strategy is used, i.e., two clusters may overlap because one paper may be related to multiple topics in a literature review. We obtain the clusters in base level as,
$C_{1}=\{c_{u}~{}|~{}c_{u}\not\subset c_{v}\},~{}u\neq v\in V_{1},~{}|c_{u}|\neq 1.$ (5)
This definition ensures that the final clusters capture distinct and relevant topics while still allowing for some overlap when a paper contributes to multiple topics. At higher levels, the clustering strategy transitions into hard clustering, as a single topic cannot belong to two different overarching topics in the taxonomy tree of the literature review. This implies that the clusters should form distinct connected components. Therefore, we generate the candidate edge set as,
=$\mathcal{E}=\{(u,v)~{}|~{}\operatorname*{arg\,max}_{v\in c_{u}}p_{uv}\},~{}u\in V_{l},~{}l>1.$ (6)
After a complete traversal of every node in $V_{l}$, $\mathcal{E}$ forms a set of connected components, resulting in disjoint clusters. This ensures that topics within the same cluster are densely connected, while maintaining clear boundaries between different topics.
Aggregation Function $g(\cdot)$. After the clustering function generates clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$, we build up the graph $G_{l+1}(V_{l+1},E_{l+1})$ at level $l+1$ by serving the clusters as the hyper-nodes and linking these nodes with the cluster density. The feature of the hyper-node is defined as the aggregation of the average feature and representative feature of the corresponding cluster:
$x_{u}=\frac{1}{|c_{l}^{i}|}\sum_{z\in c_{l}^{i}}h_{z}+h_{k},~{}\text{where}~{}k=\operatorname*{arg\,max}_{z\in c_{l}^{i}}\hat{d}_{z},$ (7)
$u\in V_{l+1}$ is the node corresponding to cluster $c_{l}^{i}$ at level $l+1$. This aggregation ensures that the hyper-node captures both the breadth (through the average) and the focus (through the distinctive feature) of the cluster. Given two hyper-nodes, $u$ and $v$, corresponding to clusters $c_{l}^{i}$ and $c_{l}^{j}$, an edge $e_{uv}\in E_{l+1}$ exists if any node in $c_{l}^{i}$ is linked to any node in $c_{l}^{j}$ at level $l$.
4.2.2 Taxonomy Tree Generation
Each cluster at every level corresponds to a specific topic within the literature review, collectively forming the taxonomy tree of the literature review. To capture relationships within and between clusters and facilitate text generation from the underlying textual graph, we employ a PLMΘ to generate the central topic for each cluster in a bottom-up manner, integrating both graph-based and text-based prompts. For a cluster $c_{l}^{i}$ at level $l$, we extract the corresponding subgraph and aggregate all relevant information into a graph embedding $h_{c_{l}^{i}}$ using $\text{GNN}_{\theta}$, then align its dimensions with the PLM’s text vector through an MLPΦ, thereby integrating the relationships between topics into the PLM. For the node $u\in V_{l+1}$ that corresponds to a cluster $c_{l}^{i}$, the topic that node $u$ represents is generated as:
$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\theta,\phi,\Phi}(y_{i}\mid y_{<i},h_{c_{l}^{i}},\{\mathcal{T}_{l}^{j}\}_{j\in c_{l}^{i}},q),$ (8)
where $q$ is the instruction to PLM, and $\mathcal{T}_{l}^{j}$ is the concatenation of the topic $Y_{l}^{j}$ and the titles of all papers under node $j\in V_{l}$. When $l=1$, $\mathcal{T}_{l}^{j}$ is simply the title of node $j$, as node $j$ at the leaf level represents a single paper. After generating the central topic for each cluster at every level, we merge the topics across different levels to form the final taxonomy tree $\hat{Y}$.
4.3 Content Generation
Once the taxonomy tree is established, we prompt the LLM, such as GPT-4o, to generate content for each topic in the taxonomy tree in parallel as $R^{i}_{l}=\text{Draft}(\tilde{Y},Y_{l}^{i},c_{l}^{i})$. The LLM is provided with the complete taxonomy tree, the specific topic of focus, and the content of relevant papers within the cluster (Appendix A.6). This ensures that the LLM not only understands the topic and references required for generation, but also the hierarchical context of the topic. After generating content for all topics, they are merged to form the complete literature review $\mathcal{R}$. Multiple versions of the literature review are generated and evaluated by the LLM, which assess aspects such as content coverage and structure. Finally, the best version of the literature review is selected as the final output.
4.4 Training Strategy
We jointly train the hierarchical clustering model and the topic generator to ensure that the hierarchical clusters formed are meaningful from a textual perspective, and that the topics generated are coherent with the structural information embedded in the citation network. The final objective function is given as,
$\operatorname*{arg\,min}_{\Theta,\theta,\phi,\Phi}\mathcal{L}=\mathcal{L}_{\text{HiCluster}}+\mathcal{L}_{\text{PLM}},$ (9)
where $\mathcal{L}_{\text{HiCluster}}$ is the loss function for hierarchical clustering, and $\mathcal{L}_{\text{PLM}}$ is the loss function for topic generation. Since only a small portion of the citation network has been labeled for hierarchical clustering (as it is challenging to collect labels that indicate which hierarchical cluster the conferences in a literature review belong to), and the learning dynamics of GNNs and LLMs differ significantly, directly training both models simultaneously can result in a complex and unstable optimization process. This leads to the GNN struggling to learn meaningful clusters early on, which subsequently hinders the LLM’s ability to generate coherent topics. Therefore, we pre-train the hierarchical clustering module to simplify the optimization process for the PLM, allowing the PLM to focus solely on content generation.
To update $\theta,\phi$. For hierarchical clustering, we train the GNN with two complementary objectives: the first optimizes clustering performance, while the second employs a hierarchical contrastive loss, inspired by Zhang et al., to ensure that nodes belonging to the same cluster at different hierarchical levels are brought closer together in the embedding space:
$\mathcal{L}_{\text{HiCluster}}=\mathcal{L}_{\text{cluster}}+$ $\mathcal{L}_{\text{HiMulCon}}=\sum_{l\in L}\frac{-1}{|E|}\sum_{(u,v)\in E}q_{uv}^{l}\log\hat{p}_{uv}^{l}+(1-q_{uv}^{l})\log(1-\hat{p}_{uv}^{l})$
$+$ $\frac{1}{|L|}\sum_{l\in L}\sum_{u\in V}\frac{-\lambda_{l}}{|S_{l}(u)|}\sum_{s_{l}\in S_{l}(u)}\log\frac{\exp(\text{sim}(h_{u},h_{s_{l}})/\tau)}{\sum_{k\in V\setminus u}\exp(\text{sim}(h_{u},h_{k})/\tau)},$ (10)
@@ 4.4 Training Strategy
where $p^{l}_{uv}$ is the probability that two nodes belong to the same cluster at level $l$. Note that this probability is always calculated for leaf-level nodes, i.e., $u,v\in V$. At higher levels ($l>1$), if $u$ and $v$ belong to two different clusters, $p^{l}_{uv}$ corresponds to the probability that these two hyper-nodes belong to the same cluster at level $l+1$. $S_{l}(u)$ represents the set of positive samples for node $u$ at level $l$, i.e., nodes that belong to the same cluster. $\lambda_{l}$ is a weighting factor for the loss at level $l$, and $\tau$ is the temperature parameter. $L$ represents the total number of hierarchical levels.
To update $\Theta,\Phi$. After pre-training $\text{GNN}_{\theta}$, we fix $\theta,\phi$ and then jointly fine-tune $\text{GNN}_{\theta}$ and $\text{PLM}_{\Theta}$ to generate the central topic for each cluster. The $\text{MLP}_{\Phi}$ maps the graph embedding $h_{c_{l}^{i}}$ in Eq. 8 to the graph token embedding $\mathbf{h}_{c}\in\mathbb{R}^{d_{\text{LLM}}}$. We leverage the text embedder function of the LLM to convert them into text embeddings $\mathbf{h}_{t}\in\mathbb{R}^{L_{t}\times d_{\text{LLM}}}$, where $L_{t}$ indicates the token length of the concatenated text contexts. Finally, the generation process in Eq. 8 becomes:.
$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\Theta,\phi,\Phi}(y_{i}\mid y_{<i},[\mathbf{h}_{c};\mathbf{h}_{t}]).$ (11)
@@
The fine-tuning of $\text{PLM}_{\Theta}$ is conducted using Low-Rank Adaptation (LoRA) (Hu et al., 2021), with $\text{MLP}_{\Phi}$ serving as an adapter for graphs, enabling efficient adjustment of both models.
5 Experiment
To demonstrate the improvement of the proposed taxonomy-then-generation method on the generation of literature review papers, we conducted comprehensive experiments. In addition to analyzing the generative quality of the literature review, this section will answer the following questions:
  • RQ1. Why use hierarchical clustering instead of a simple clustering strategy?
  • RQ2. Why do we need a taxonomy tree to guide the generation process?
  • RQ3. Why not prompt LLMs to generate taxonomy trees and then literature reviews?
Setup. The fine-tuned topic generator is LLaMA (Touvron et al., 2023), while GPT-4o serves as the content generator [^3] Specifically, we use LLaMA-2-7b and gpt-4o-2024-05-13.. For hierarchical clustering, we employ the GAT (Veličković et al., 2017) in hierarchical clustering. Implementation details can be found in Appendix A.2.
5.1 Dataset
@@
We manually collected 518 high-quality literature review articles with clear taxonomy or well-defined article structures from arXiv [^4] https://arxiv.org/. Most of these review papers were published within the past three years. For each literature review, we extracted its taxonomy tree and gathered its 2-hop citation network, which includes the direct citations of the review paper and the citations of the review’s references. The mutual citation relationships among these references form a complex citation network for each literature review. These 2-hop citation networks contain an average of 6,658.4 papers and 11,632.9 edges, accommodating isolated papers. More details are in Appendix A.1.
5.2 Evaluation
We compare the reviews generated by our model with those written by human experts, zero-shot LLMs and naive RAG-based LLMs, i.e., GPT-4o and Claude-3.5. While zero-shot LLMs rely solely on their pretrained knowledge to generate literature review content, naive RAG-based LLMs utilize a simple BM25 retriever. The LLM backbone of AutoSurvey and HiReview are both GPT-4o. Additionally, we benchmarked our model against the state-of-the-art review generator, AutoSurvey (Wang et al., 2024). We evaluate the quality of generated content with LLMScore and BertScore (Zhang et al., 2019). For LLMScore, Wang et al. demonstrated that LLM-based evaluations of literature review align well with human preferences. Therefore, we employed multiple LLMs to assess the overall quality of the generated content and averaged their evaluation scores. Following AutoSurvey’s scoring criteria, we evaluate the generated reviews based on coverage, structure, and relevance when selecting the best output and calculating the LLMScore. When evaluate the review content, instead of directly scoring the generation, we have LLMs compare the generated content with human-written reviews (Appendix A.6).

### v2

@@ 4.2. Hierarchical Citation Graph Clustering
where $\mathcal{N}(u)$ indicates the neighboring nodes of node $u$. The density assesses how densely connected each node is within its local neighborhood. High-density nodes are more likely to be in the core of a cluster, whereas low-density nodes are more likely to be in ambiguous regions between clusters. $\hat{d}_{u}=0$ if $u$ is isolated.
To balance fine-grained topic relationships with a well-structured taxonomy, we employ soft clustering at the first level, allowing topic overlap to reflect the interdisciplinary nature of research, and progressively transition to hard clustering at higher levels to enforce distinct, non-overlapping categories in the taxonomy hierarchy.
At the base level ($l=1$), given $\hat{p}^{(l)}_{uv}$, $\hat{d}_{u}$, $\hat{d}_{v}$, and a pre-defined edge connection threshold $p_{\tau}$, we generate the candidate cluster as:
(4) $c^{(l)}_{i}=\{u\}\cup\{v\mid\hat{d}_{u}<\hat{d}_{v}\ \text{and}\ \hat{p}^{(l)}_{uv}>p_{\tau}\}.$
Nodes with higher density may belong to multiple clusters, reflecting the interdisciplinary nature of research. Final clusters are formed by merging candidate clusters:
(5) $\mathcal{C}^{(l)}=\{c^{(l)}_{i}\mid c^{(l)}_{i}\not\subset c^{(l)}_{j},\ |c^{(l)}_{i}|\neq 1\}.$
At higher levels ($l>1$), clustering transitions to hard clustering to enforce distinct, non-overlapping categories. Edges between nodes are formed based on the strongest connection probabilities:
(6) =$\mathcal{E}=\{(u,v)\mid\arg\max_{v\neq u\in V}\hat{p}^{(l)}_{uv}\},\quad u\in V^{(l)}.$
After traversing all nodes, connected components defined by $\mathcal{E}$ form disjoint clusters $\mathcal{C}^{(l)}$, ensuring clear topic boundaries within the taxonomy hierarchy.
4.2.2. Aggregation Operator $\text{AGG}(\cdot)$.
The aggregation function computes the feature representation for each hyper-node in $V^{(l+1)}$. Specifically, the feature of a hyper-node is derived from both the average and the most representative node features within its corresponding cluster:
(7) $x_{u}=\frac{1}{|c^{(l)}_{i}|}\sum_{z\in c^{(l)}_{i}}h_{z}+h_{k},\quad\text{where}\quad k=\arg\max_{z\in c^{(l)}_{i}}\hat{d}_{z}.$
Here, $h_{k}$ corresponds to the node with the highest density in cluster $c^{(l)}_{i}$, ensuring that the hyper-node captures both general and distinctive characteristics of the cluster. This aggregated feature set $X^{(l+1)}$ serves as the input for the next hierarchical level.
4.2.3. Hierarchical Citation Graph Clustering Objective.
Since paper-level annotations are the most granular and reliable supervision, leveraging them indirectly optimizes higher-level clusters, preserving semantic integrity and aligning the taxonomy with real-world research structures without relying on hypothetical hypernode labels. Moreover, papers consistently clustered together across multiple hierarchical levels should be closer in the embedding space than those grouped only at lower levels or not at all, reflecting stronger and more persistent semantic relationships. Therefore, we design the objective function to operate on paper-level labels as:
(8) $\mathcal{L}_{\text{HiCluster}}=\mathcal{L}_{\text{cluster}}+\alpha\cdot\mathcal{L}_{\text{HiMulCon}},$
where the first term, $\mathcal{L}_{\text{cluster}}$, optimizes clustering by grouping papers with structural and semantic similarities at each hierarchical level. The second term, $\mathcal{L}_{\text{HiMulCon}}$, is a hierarchical contrastive loss that enforces multi-level consistency by bringing together nodes that persist in the same cluster across levels and pushing apart those that diverge, ensuring the embeddings reflect the taxonomy’s hierarchical structure. $\alpha$ controls the contribution from the contrastive loss. Specifically, the clustering loss term is defined as follows,
(9) $\mathcal{L}_{\text{cluster}}=\sum_{l=1}^{L}\frac{-1}{|E|}\sum_{(u,v)\in E}q_{uv}^{(l)}\log\hat{p}_{uv}^{(l)}+(1-q_{uv}^{(l)})\log(1-\hat{p}_{uv}^{(l)})$
where $q_{uv}^{(l)}=1$ if $u,v\in V$ belong to the same cluster at level $l$, and $q_{uv}^{(l)}=0$ otherwise. This is computed over the edge set $E$ rather than all node pairs, as we empirically found it reduces computational costs without significantly affecting performance.
To ensure that papers consistently assigned to the same cluster across multiple levels are positioned closer in the embedding space, the second term is formulated as a hierarchical contrastive loss as:
(10) $\mathcal{L}_{\text{HiMulCon}}=\frac{1}{|L|}\sum_{l=1}^{L}\sum_{u\in V}\frac{-\delta_{l}}{|S^{(l)}_{u}|}\sum_{k\in S^{(l)}_{u}}\log\frac{\exp(\text{sim}(h_{u},h_{k})/\tau)}{\sum_{v\in V\setminus u}\exp(\text{sim}(h_{u},h_{v})/\tau)}$
@@
where $S^{(l)}_{u}$ represents the set of positive samples for node $u$ at level $l$, i.e., nodes that belong to the same cluster. $\delta_{l}$ is a weighting factor that reflects the importance of each level for the loss, and $\tau$ is the temperature parameter. $L$ represents the total number of hierarchical levels.
4.3. Hierarchical Taxonomy Node Verbalization
A citation graph $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$ reveals local research communities that expand into larger clusters through hierarchical analysis. Its structure supports a bottom-up taxonomy, where denser subgraphs form the basis for broader categories. After hierarchical citation graph clustering, each cluster $c^{(l)}\in\mathcal{C}^{(l)}$ corresponds to a topic $u\in U$ in the taxonomy $\mathcal{T}$. We now introduce an iterative generation method to achieve the mapping $h:\{D_{v}\}_{v\in V}\to\{D_{u}\}_{u\in U}$ for summarizing the central concept of each cluster. Specifically, we propose an iterative graph-taxonomy generation strategy that refines fine-grained topics and abstracts them into higher-level concepts, ensuring semantic consistency across hierarchy levels.
4.3.1. Iterative Generation
Given the hierarchical decomposition $\mathcal{G}^{(l)}=(V^{(l)},E^{(l)},X^{(l)})$ of a citation graph $\mathcal{G}$ at level $l$ and its corresponding clusters $\mathcal{C}^{(l)}$, we generate the central concept for each cluster. To generate a meaningful taxonomy for a specific topic $q$, we employ a pre-trained LLMΘ as the backbone for concept abstraction, leveraging its extensive text processing capabilities. For each cluster $c^{(l)}\in\mathcal{C}^{(l)}$, the probability distribution of its generated central concept is defined as:
(11) $p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q)=\prod_{i=1}^{r}p_{\Theta,\Phi}(y_{i}\mid y_{<i},c^{(l)},q),$
where $Y_{u}^{(l+1)}$ denotes the token sequence of central concept of cluster $c^{(l)}\in\mathcal{C}^{(l)}$, corresponding to the hyper-node $u\in V^{(l+1)}$. $q$ represents the instruction prompt reflecting user preferences, and $y_{<i}$ denotes the prefix tokens. $\Phi$ indicates the parameters of a projector that maps the graph embedding into the LLM text embedding space.
Each cluster at level $l$ represents a subgraph of $\mathcal{G}^{(l)}$. To capture interactions between hyper-nodes for central concept generation, we apply a projector MLPΦ to map graph embeddings into the LLM text space as graph tokens, where graph embeddings are produced using the same GNNθ in the hierarchical citation graph clustering process. While hierarchical aggregation retains cluster features in hyper-nodes, information loss from individual papers is inevitable as the hierarchy deepens. To address this, we also incorporate paper-level information from the base graph, as each hyper-node at higher levels can be coarsened to its corresponding set of base-level papers. Formally, the cluster $c^{(l)}$ is encoded as:
(12) $\mathbf{h}_{c}=\text{MLP}_{\Theta}(\text{GNN}_{\theta}(c^{(l)}))\in\mathbb{R}^{d_{\text{LLM}}}$
where $d_{\text{LLM}}$ denotes the LLM text embedding dimension. The associated papers and instructions are encoded using the LLM text embedder as:
(13) $\mathbf{h}_{\pi}=\text{TextEmbedder}(\{D_{v}\}_{v\in\pi_{c}^{(l)}})\in\mathbb{R}^{N_{\pi}\times d_{\text{LLM}}},$
(14) $\mathbf{h}_{q}=\text{TextEmbedder}(q)\in\mathbb{R}^{N_{q}\times d_{\text{LLM}}},$
where $N_{\pi}$, $N_{q}$ indicate the length of token sequences, $\pi_{c}^{(l)}$ represents the set of base-level papers corresponding to the coarsened $c^{(l)}$. Finally, the probability distribution in Equation (11) is as,
(15) $p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q)=\prod_{i=1}^{r}p_{\Theta,\Phi}(y_{i}\mid y_{<i},[\mathbf{h}_{c};\mathbf{h}_{\pi};\mathbf{h}_{q}]),$
where [;][~{};~{}][ ; ] indicates the concatenation of token embeddings before feeding them through transformer layers of LLMΘ.
4.3.2. Hierarchical Generation Objective
The goal of taxonomy generation is to maximize the likelihood of generating coherent central concepts for all clusters across the hierarchical citation graph, thereby constructing the entire taxonomy tree. Formally, the objective is defined as:
(16) $\mathcal{L}_{\text{Gen}}=\sum_{l=1}^{L}\sum_{c^{(l)}\in\mathcal{C}^{(l)}}\log p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q).$
By maximizing this hierarchical likelihood, the model generates semantically consistent and contextually relevant topics, ensuring coherence throughout the taxonomy tree.
4.4. Optimization for CGT
We jointly train the hierarchical citation graph clustering model and the taxonomy node verbalizer to ensure that the hierarchical clusters formed are meaningful from a textual perspective, and that the topics generated are coherent with the structural information embedded in the citation network. The final objective function is given as,
(17) $\operatorname*{min}_{\Theta,\Phi,\theta,\phi}\mathcal{L}=\mathcal{L}_{\text{Gen}}(\Theta,\Phi)+\lambda\cdot\mathcal{L}_{\text{HiCluster}}(\theta,\phi),$
where $\lambda$ is a pre-defined parameter, $\mathcal{L}_{\text{HiCluster}}$ is the loss function for hierarchical citation graph clustering, and $\mathcal{L}_{\text{Gen}}$ is the loss function for taxonomy node verbalization. Since only a small portion of the citation graph has been labeled for hierarchical citation graphclustering (as it is challenging to collect labels that indicate which hierarchical citation graph cluster the conferences in a literature review belong to), and the learning dynamics of GNNs and LLMs differ significantly, directly training both models simultaneously can result in a complex and unstable optimization process. This leads to the GNN struggling to learn meaningful clusters early on, which subsequently hinders the LLM’s ability to generate coherent topics. Therefore, we pre-train the hierarchical citation graph clustering module to simplify the optimization process for the LLM, allowing the LLM to focus solely on content generation.
@@ 4.4. Optimization for CGT
After pre-training the hierarchical clustering module $(\theta,\phi)$, we then jointly fine-tune them and the concept generator $(\Theta,\Phi)$ to generate the central topic for each cluster. The fine-tuning is conducted using Low-Rank Adaptation (LoRA) (Hu et al., 2021), with $\text{MLP}_{\Phi}$ serving as an adapter for graphs, enabling efficient adjustment of both models.
5. Experiment
To demonstrate the effectiveness of the proposed taxonomy learning method, we evaluate the quality of the generated taxonomies and further assess the model’s ability to guide literature review generation. We first introduce the experimental settings in Section 5.1. Next, we present the performance of HiGTL on citation graphs in the computer science domain in Section 5.2. We then conduct ablation studies to demonstrate the contributions of Hierarchical Citation Graph Clustering and Taxonomy Node Verbalization in Section 5.3. Finally, we include a case study to highlight the strengths and limitations of our framework in Section 5.4.
5.1. Experiment Setting
5.1.1. Datasets
@@
We manually collected 518 high-quality literature review articles with clear taxonomies or well-defined structures from arXiv [^1] https://arxiv.org/ across various computer science domains, with most published in the past three years. To evaluate taxonomy learning performance, we extracted taxonomy trees from these reviews and gathered the 1-hop citation graph for each review, representing citation relationships among its direct references. Additionally, we collected the 3-hop citation graph for each review to assess literature review generation based on HiGTL-generated taxonomies. Further details are provided in Appendix A.1.
5.1.2. Comparison methods
We compare our model with the following state-of-the-art taxonomy learning methods: HiExpan (Shen et al., 2018b), which constructs taxonomies by recursively expanding seed sets using pattern-based and distributional embedding methods, followed by a global optimization module to refine the taxonomy structure for better coherence. TaxoGen (Zhang et al., 2018), which generates topic taxonomies by embedding concept terms into a latent space and applying adaptive spherical clustering and local term embeddings to ensure fine-grained semantic distinctions at each hierarchical level. NetTaxo (Shang et al., 2020), which integrates text and network data for taxonomy construction by learning term embeddings from both contexts and employing instance-level motif selection to enhance hierarchical citation graph clustering and topic differentiation. Since our model leverages LLMs, we also include GPT-4o and Claude-3.5 as baselines for taxonomy generation. Implementation details are provided in Appendix A.2.
5.1.3. Evaluation Metrics
To comprehensively evaluate the model’s generation quality, we assess outputs from three perspectives: LLMScore, Human Evaluation, and BertScore (Zhang et al., 2019). Wang et al. demonstrated that LLM-based evaluations align closely with human preferences, so we leverage LLMs to rate the generated content. We also employ BertScore to measure the semantic similarity between generated texts and human-written references. However, human evaluation remains essential for accurately capturing human preferences and for cases without available human-written labels, ensuring a thorough assessment of the generated content, especially given LLM hallucinations. Details of the prompts used for LLM evaluation and the instructions for the human evaluation process are provided in Appendix A.3.
5.2. Main Results
5.2.1. Taxonomy Learning Results
Results in Table 1 demonstrate that HiGTL outperforms all other models across multiple evaluation metrics, achieving the highest scores in Coverage (0.9357), Structure (0.9413), Relevance (0.8748), and the overall LLM Average (0.9173). It also leads in Human Evaluation metrics with the highest Adequacy (0.7150) and Validity (2.6700), reflecting its superior quality in both automated and human assessments. Additionally, HiGTL achieves the highest BertScore (0.8694), indicating strong alignment with reference taxonomies. Among large language models, GPT-4o shows competitive performance, particularly in Relevance (0.8828), where it slightly edges out HiGTL. However, it falls behind in other areas like Structure and Coverage, as well as human evaluations. Traditional taxonomy learning methods like NetTaxo and TaxoGen consistently lag behind HiGTL and LLM-based approaches due to their lack of ability to capture continuous, nuanced concepts.